Prescott Lumber processes logs into grade A and grade B lumber. Logs cost $20,200 per load. The milling process produces 3,000 units of grade A with a market value of $58,800, and 12,000 units of grade B with a market value of $8,400. The cost of the milling process is $7,000 per load.

Required: a. If the costs of

Prescott Lumber processes logs into grade A and grade B lumber. Logs cost $20,200 per load. The milling process produces 3,000 units of grade A with a market value of $58,800, and 12,000 units of grade B with a market value of $8,400. The cost of the milling process is $7,000 per load.

Required:

a. If the costs of the logs and the milling process are allocated on the basis of units of output, what cost will be assigned to each product? (Do not round intermediate calculations.)

Cost Allocation
Grade A

Grade B

b. If the costs of the logs and the milling process are allocated on the basis of the net realizable value, what cost will be assigned to each product? (Do not round intermediate calculations.)

Cost Allocation
Grade A
Grade B

c-1. How much profit or loss does the grade B lumber provide using the data in this problem and your analysis in requirement (a)? (Do not round intermediate calculations.)
Answer #1

a.Calculation of Cost allocation of each product on the basis of units of output

Grade A = proportionate logs cost and cost of milling.

= (20200+7000)*3000/15000=$5440

Grade B = 27200*12000/15000=$21760

b.Calculation of cost allocation of each product on the basis of net realizable value

Grade A = 27200*58800/67200= $23800

Grade B = 27200*8400/67200 = $3400

c-1 Calculation of profit/ loss if Grade B lumber in requirement(a)

Sales - cost = 8400-21760=$13360(Loss)

​​​​​